OpenStudy (cwrw238):

I'm having trouble solving this differential equation (1 - x^2) y' - xy = x I'm not sure about the integrating factor is it (1 - x^2) ^(1/2) or 1/ ((1 - x^2)^(1/2) (or maybe its neither!). Please help.

OpenStudy (phi):

if in the form y' + P(x) y = Q(x) the integrating factor is e^F where F= integral P(x) dx

OpenStudy (anonymous):

\[(1-x^2)\frac{dy}{dx} -xy = x\]

OpenStudy (cwrw238):

@ph1 - yes - I remembered that I'm (to my shame) confuse by the algebra

OpenStudy (phi):

integral -x dx/(1-x^2) = 1/2 * ln(1-x^2) (toss in abs signs to be sure it's positive)

OpenStudy (phi):

exp (ln(1-x^2)^(1/2))= sqrt(1-x^2)

OpenStudy (phi):

integral du/u = ln(u) in this case u = 1-x^2 du = -2x dx

OpenStudy (phi):

\[ e^{ab} = \left(e^a\right)^b \] so \[ e^{\ln(1-x^2)\cdot \frac{1}{2}}=\left(e^{\ln(1-x^2)}\right)^\frac{1}{2} \] \[= (1-x^2)^\frac{1}{2}= \sqrt{1-x^2} \]

OpenStudy (cwrw238):

ok - i see it now . I was confused with the sign when integrating thanx

OpenStudy (cwrw238):

i need more practice in FODEs .......

OpenStudy (phi):

are you saying you can't finish ?

OpenStudy (cwrw238):

no i think i've got it now I get the solution to be y = A(1 - x^2)^(-1/2) - 1

OpenStudy (phi):

that looks good

OpenStudy (cwrw238):

I got it down to